OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). 7 tan3x - 21 tan x = 0

zepdrix (zepdrix):

Is that \(\large\rm \tan(3x)\) or \(\large\rm \tan^3x\) ? :)

OpenStudy (anonymous):

tan ^3 x

zepdrix (zepdrix):

\[\large\rm 7\tan^3x-21\tan x=0\]Let's factor some stuff first. Looks like they both have..... a tangent... and a 7, ya?

OpenStudy (anonymous):

7tanx(tan^2x-3)=0

zepdrix (zepdrix):

Ok good! Now we apply our `Zero-Factor Property`, setting each individual factor equal to zero, and then solving from there. So our first factor equal to zero is: \(\large\rm 7\tan x=0\)

zepdrix (zepdrix):

You can divide by 7 to simplify things down a bit: \(\large\rm \tan x=0\) Understand how to solve for x in this case? :)

OpenStudy (anonymous):

would you plug it into a calculator using tan^-1?

zepdrix (zepdrix):

I guess you `could` do that :p it's better just remember some of your special angles. tangent is 0 when the angle x is 0.

zepdrix (zepdrix):

Oh but I guess umm... they want all of the solutions from 0 to 2pi, so that produces another angle as well.

OpenStudy (anonymous):

oh yeahhhh, so then pi and 0? or would it be 1/pi and 3/pi?

zepdrix (zepdrix):

sorry got distracted :) um um... yaaaaa, 0 and pi sound good for our first two solutions.

zepdrix (zepdrix):

Applying the zero-factor property again:\[\large\rm \tan^2x-3=0\]

zepdrix (zepdrix):

So ummm... how bout.. add 3. then square root or something, ya?

zepdrix (zepdrix):

\[\large\rm \tan^2=3\]

zepdrix (zepdrix):

\[\large\rm \tan x=\pm\sqrt{3}\]

zepdrix (zepdrix):

The positive root 3 will give you 2 angles again. while the negative root 3 will give you 2 more angles! So this solution set is actually giving us 4 angles! kinda crazy.

OpenStudy (anonymous):

uhhhhh... how do you get your solutions from \[\pm \sqrt{3}\]

zepdrix (zepdrix):

Hmm you gotta get better with your unit circle missy! :O http://4.bp.blogspot.com/-FMVojhlkcSQ/UInlIKO88oI/AAAAAAAAAC4/-fzMOz6di2Y/s1600/image010.jpg So umm.. i couldn't find a really good picture, but here is one.

zepdrix (zepdrix):

The coordinate pair lists first the sine value of that angle, and then the cosine. And then outside of the brackets, on the right, is the tangent value of that angle. So if we look in the first quadrant, it looks like the angle pi/3 gives us sqrt(3), ya?

OpenStudy (anonymous):

Ohhhhh! I see now! :P Thank you so much!

zepdrix (zepdrix):

cool c: